# Taming Modality Entanglement in Continual Audio-Visual Segmentation

### v1

@@ Abstract
Recently, significant progress has been made in multi-modal continual learning, aiming to learn new tasks sequentially in multi-modal settings while preserving performance on previously learned ones. However, existing methods mainly focus on coarse-grained tasks, with limitations in addressing modality entanglement in fine-grained continual learning settings. To bridge this gap, we introduce a novel Continual Audio-Visual Segmentation (CAVS) task, aiming to continuously segment new classes guided by audio. Through comprehensive analysis, two critical challenges are identified: 1) multi-modal semantic drift, where a sounding objects is labeled as background in sequential tasks; 2) co-occurrence confusion, where frequent co-occurring classes tend to be confused. In this work, a Collision-based Multi-modal Rehearsal (CMR) framework is designed to address these challenges. Specifically, for multi-modal semantic drift, a Multi-modal Sample Selection (MSS) strategy is proposed to select samples with high modal consistency for rehearsal. Meanwhile, for co-occurence confusion, a Collision-based Sample Rehearsal (CSR) mechanism is designed, allowing for the increase of rehearsal sample frequency of those confusable classes during training process. Moreover, we construct three audio-visual incremental scenarios to verify effectiveness of our method. Comprehensive experiments demonstrate that our method significantly outperforms single-modal continual learning methods.
I Introduction
Humans are inherently capable of continuously learning while retaining knowledge from previous tasks. For example, infants can progressively recognize new animals while remembering those they have already learned. This human ability has motivated extensive research into continual learning [1], which enables models to learn sequential tasks. Early work [2, 3, 4] primarily focused on classification, employing techniques such as regularization or rehearsal to mitigate catastrophic forgetting. Subsequent methods [5] have extended continual learning to semantic segmentation. However, when directly applied to multi-modal (e.g. audio-visual) scenarios, these single-modal methods exhibit suboptimal performance [6].
@@ I Introduction
Recently, several methods [7, 8, 6] have extended continual learning to multi-modal scenarios. For example, AV-CIL [7] proposes a continual audio-visual classification method with a dual similarity constraint enforcing both instance-level and class-level cross-modal semantic consistency. ContAV-Sep [8] proposes a framework for audio-visual separation that incorporates cross-modal similarity distillation to preserve semantic consistency between modalities. Meanwhile, real-world applications require fine-grained audio-visual continual learning. For example, embodied intelligence needs to identify the source of a vocalization from environmental audio-visual cues. However, existing methods primarily focus on coarse-grained audio-visual tasks and therefore fail to address fine-grained tasks, such as disentangling pixel-level visual features from audio signals under continual learning scenarios.
Meanwhile, recent research [9, 10, 11] has explored fine-grained modality entanglement between audio signals and visual features in audio-visual segmentation. AVSBench [9] establishes the first benchmark for aligning the pixel-level visual semantics with the corresponding audio signals. COMBO [11] further explores bilateral relations of three entanglements, pixel, modality, and temporal, to enhance the model’s representational capacity. However, audio-visual segmentation cannot be directly applied to continual learning scenarios, as it is designed for static settings.
To this end, we introduce a novel fine-grained multi-modal continual learning task, termed Continual Audio-Visual Segmentation (CAVS). Specifically, CAVS needs to perform audio-visual segmentation in a sequential task setting while retaining knowledge of previously seen classes. To address CAVS, we reformulate the AVS [9, 10] framework and adapt classical continual semantic segmentation methods to the audio-visual context. Based on our observations, we identify two new challenges in fine-grained continual learning tasks: (1) Multi-modal semantic drift: Incorrect audio-visual semantic alignment (e.g. drum-background) due to mislabeling of learned classes as background exacerbates catastrophic forgetting. (2) Co-occurrence confusion: Frequent co-occurrence of categories leads to modality entanglement, for example, the audio modality of woman becomes entangled with visual modality of guitar in Fig. 1. In essence, these two issues are manifestations of modality entanglement from different perspectives.
@@
To validate the effectiveness of CMR, we reformulate the audio-visual dataset AVSBench [9] into three sequential task setup to better simulate a continual learning scenario. Specifically, our datasets include (1) AVSBench-Class Incremental (AVSBench-CI), (2) AVSBench-Class Incremental for Single-object (AVSBench-CIS), and (3) AVSBench-Class Incremental for Multi-object (AVSBench-CIM). Comprehensive experiments demonstrate that our proposed method achieves encouraging performance, showcasing its ability to effectively address the multi-modal semantic drift and co-occurrence confusion in CAVS.
@@
II Related Work
@@
II-A Continual Learning.
Continual learning focuses on incrementally training models to adapt to new tasks while preserving knowledge from previously learned ones. Recently, many works  [4, 3, 12, 2, 13, 14, 15, 16, 17] have proposed rehearsal-based and rehearsal-based methods to address the problem of catastrophic forgetting. Rehearsal-based methods [4, 3, 12, 2, 13] allow for the storage of a small subset of old data in memory, which is later utilized for rehearsal during training. iCaRL [2] introduces a strategy to identify and retain the most representative samples for each class, which are replayed during training to mitigate forgetting in class-incremental learning. Pseudo-sample rehearsal-based methods [14, 15, 16] utilize generative models to create pseudo-samples of old classes. DGR [15] establishes an initial framework where learning each new task is coupled with replaying the data generated by the old generative model. Building upon continual learning [2, 13, 12, 14, 15, 16], Class-Incremental Semantic Segmentation (CISS) requires pixel-level classification to achieve fine-grained segmentation [5, 18, 19, 20, 21, 22]. PLOP [18] suggests generating pseudo-labels by identifying latent past classes within the current background. ScaleSeg [20] employs prototypes refined through online contrastive clustering and incorporates a background diversity strategy to boost plasticity. While Cermelli et al. (2020) addressed semantic shifts within a single modality, our work reveals more complex multi-modal semantic drift where modal consistency is considered.
II-B Audio Visual Segmentation
Audio-visual segmentation (AVS) is a novel and challenging task that localizes sound sources in visual scenes by pixel-level prediction [9, 10, 11, 23, 24, 25, 26, 27, 28]. AVSBench [9] establishes the first audio-visual segmentation benchmark and introduces the Temporal Pixel-wise Audio-Visual Interaction (TPAVI) module to incorporate audio semantics as guidance for visual segmentation. AVSegFormer [10] develops a transformer-based framework with audio queries, learnable queries, and an audio-visual mixer for selective attention and dynamic feature adjustment. CATR  [24] proposes a combinatorial fusion framework that captures audio-visual spatiotemporal dependencies through cross-modal interaction modelling. ECMVAE [23] decomposes audio and visual data in latent space, explicitly modeling both shared and modality-specific representations to enhance segmentation performance. COMBO [11] rethinks AVS by exploring the bilateral relations of three entanglements, pixel, modality, and temporal, to enhance the model’s representation ability. In this work, we develop a framework for continual learning scenarios, making the task more aligned with real-world applications.
II-C Multimodal Learning
Multimodal learning [29, 30, 31, 32] focuses on integrating information across diverse modalities and investigating the intricate interrelationships between them in various contexts. Wei [29] first estimates each modality’s learning status based on separability in its unimodal representation space, then uses this to softly initialize the corresponding unimodal encoder. MMPareto [30] employs gradient-based optimization to mitigate model bias towards specific modalities during training, thereby enhancing multimodal learning performance. MMH [33] proposes a multimodal reconstruction framework that guides the reconstruction network to directly learn modality-shared representations from the multimodal encoder, thereby capturing richer cross-modal interactions. To compared with more recent work, Finger [31] focuses ondistinguishing foreground from background and transfer-ring unimodal knowledge, while we focus on selecting con-sistent samples through modal contribution and replaying them according to collision frequency. From task level, Finger aims to seamlessly integrate new classes with limitedincremental samples, while we focus on avoiding interfer-ence with old task knowledge when training on new tasks. Meanwhile, in contrast to Open-set AVS, continual learning AVS deals with learning from a continuous data stream under memory constraints, without revisiting past data.
III Method
@@
III-A Revisiting Continual Semantic Segmentation (CSS)
CSS assumes that tasks arrive sequentially, with each task containing a set of categories $\mathcal{C}^{t}$ and a corresponding training set $\mathcal{D}_{t}$, where $t$ denotes the current learning stage. The goal of the learning task $\mathcal{D}_{t}$ at a given stage $t$ is to learn a model $f_{\theta}^{t}$ parameterized by $\theta^{t}$ to accurately predict the label given an input image $X$. The predicted output segmentation mask for pixel $i$ can be computed as:
$y_{i}=\arg\max\{f_{\theta^{t}}(X)[i,c]\}^{|\mathcal{Y}|-1}_{c=0},$ (1)
where $f_{\theta^{t}}(X)[i,c]$ denotes the predicted probability of class $c$ at pixel $i$.
@@ III-A Revisiting Continual Semantic Segmentation (CSS)
III-B Problem Setup and Notation of CAVS
For CAVS, the input space is defined as $\mathcal{S}\subset\mathcal{X}\times\mathcal{A}$, where $\mathcal{X}$ and $\mathcal{A}$ represent the visual and audio modalities, respectively. Each input sample $S=(\{S^{k}_{v}\},S_{a})\in\mathcal{D}_{t}$ contains $T$ consecutive video frames paired with an audio signal $S_{a}$, where $T=10$. The sounding objects in the $k$-th video frame $S^{k}_{v}$ are annotated with pixel-level labels. The objective of the $t$-th learning stage is to learn a model $f_{\theta^{t}}^{v,a}:\mathcal{S}\mapsto\mathbb{R}^{N\times|{\mathcal{C}^{t}}|}$, where $N$ is the number of pixels per frame. In this setting, the segmentation mask for pixel $i$ can be computed as follows:
$y_{i}=\arg\max\{f_{\theta^{t}}^{v,a}(\{S_{v}^{k}\},S_{a})[i,c]\}^{|\mathcal{Y}|-1}_{c=0}.$ (2)
In contrast, for task $\mathcal{D}_{t}$, both non-sounding objects from $\mathcal{Y}^{t}$ and sounding objects from $\mathcal{Y}^{t-1}$ are assigned the background label, while the audio $S_{a}$ remains unchanged. Compared to AV-ICL [7], CAVS demands more substantial fine-grained alignment between global audio cues and local visual semantics.
Figure 3: Illustration of inter-modal consistent samples and entanglement of modality. (a) Cases 1 and 4 don’t appear in practice because selection uses already well-trained samples where audio and video predictions match the ground truth. Case 3 represents samples characterized by multi-modal semantic drift and is typically excluded due to substantially large $|\Delta(S_{a})|$. Conversely, Case 2 is kept because of its cross-modal semantic consistency. (b) Classes with infrequent co-occurrence exhibit weak audio-visual entanglement, while frequent co-occurrence leads to strong cross-modal entanglement (e.g., guitar sounds and images of women).
III-C Multi-modal Sample Selection
@@
Inspired by the work in [34], where Shapley values are leveraged to quantify uni-modal contributions to model predictions, we propose a Multi-modal Sample Selection (MSS) strategy. By quantifying the contribution of the audio modality, this strategy identifies samples with high inter-modal consistency for rehearsal. Formally, given a video sample $S=(\{S^{k}_{v}\},S_{a})\in\mathcal{D}_{t}$, we train two parallel models:
$f_{\theta_{t}}^{v}(\{S^{k}_{v}\}):\mathcal{X}\mapsto\mathbb{R}^{N\times|{\mathcal{Y}^{t}}|},$ (3)
$f_{\theta_{t}}^{v,a}(\{S^{k}_{v}\},S_{a}):\mathcal{S}\mapsto\mathbb{R}^{N\times|{\mathcal{Y}^{t}}|}.$ (4)
@@
For each newly added class $c\in\mathcal{C}^{t}$, we select the top-$k$ samples with the smallest absolute audio contribution deviation $|\Delta(S_{a})|$ from $\mathcal{D}_{t}$ to construct the memory buffer $M_{t}$.
These selected samples are stored and replayed during the training of subsequent tasks through $\mathcal{D}_{t+1}\cup M_{t}$, which effectively reinforces cross-modal associations. Our ablation studies demonstrate that this criterion outperforms random selection by 2.0 mIoU (see Tab. IV), highlighting the importance of modality consistency in sample rehearsal.
III-D Collision-based Sample Rehearsal
@@
To implement this idea, we propose the Collision-based Sample Rehearsal (CSR) strategy, which identifies classes prone to co-occurrence confusion by detecting collisions between the old model’s predictions and the ground truth. As illustrated in Fig. 4, for a new sample $S$, a collision occurs when the old model $f^{v,a}_{\theta_{t-1}}$ predicts an old class $c_{old}\in\mathcal{Y}^{t-1}$ in a spatial position where the ground truth $c_{new}\in\mathcal{C}^{t}$ appears. Specifically, with the old model and task $\mathcal{D}_{t}$, the collisions between the prediction of $f^{v,a}_{\theta_{t-1}}$ and $\mathcal{D}_{t}$ is first computed. Inferring the video $S$ with the old model $f^{v,a}_{\theta_{t-1}}$, we obtain a collision pair $(c_{old},c_{new})$. Since the old model has not trained on new samples, it can only predict old classes $c_{old}\subset\mathcal{Y}^{t-1}$. Assuming that the predicted result is $c_{old}$ and the ground truth label is $c_{gt}$, we count all collision pairs $(c_{old},c_{new})$ and identify the learned class with the highest number of collisions as the most confusing class for the current video $S$:
$\mathcal{P}(S)=\arg\max\{Count(c_{i},c_{j})|i\in\mathcal{Y}^{t-1},j\in\mathcal{C}^{t}\}.$ (8)
@@
$R_{c}=\frac{Count(\mathcal{P}(S)=c)}{\sum\{Count(c_{i},c_{j})|i\in\mathcal{Y}^{t-1},j\in\mathcal{C}^{t}\}},$ (9)
@@
Algorithm 1 Collision-based Sample Rehearsal
0: Old model $f^{v,a}_{\theta^{t-1}}$, Training dataset $\mathcal{D}_{t}$, Semantic label $\mathcal{Y}_{gt}$, Threshold $\mathcal{T}$
0: Collision frequency $\mathcal{F}$
1:for $S_{i}\in\mathcal{D}_{t}$ do
2:  Compute $\hat{\mathcal{Y}}\leftarrow f^{v,a}_{\theta^{t-1}}(S_{i})$
3:  Mask $\mathcal{M}\leftarrow(\hat{\mathcal{Y}}\neq background)\land(\mathcal{Y}\neq background)$
4:  Collision Region $\mathcal{I}\leftarrow(\hat{\mathcal{Y}}\neq\mathcal{Y})\land\mathcal{M}$
5:   $\triangleright$ Count Pairs:
6:  for $i\in\mathcal{I}$ do
7:   $Collision(\hat{\mathcal{Y}}_{i},\mathcal{Y}_{i})=Collision(\hat{\mathcal{Y}}_{i},\mathcal{Y}_{i})+1$
8:  end for
9:   $\triangleright$ Get Most Confused Class:
10:  $(\mathcal{C}_{old},\mathcal{C}_{new})\leftarrow\arg\max(Collision(\hat{\mathcal{Y}_{\mathcal{I}}},\mathcal{Y}_{\mathcal{I}})$
11:  $\mathcal{R}\leftarrow\frac{Collision(\mathcal{C}_{old},\mathcal{C}_{new})}{\sum Collision(\hat{\mathcal{Y}_{\mathcal{I}}},\mathcal{Y}_{\mathcal{I}})}$
12:   $\triangleright$ Update Frequency:
13:  if $\mathcal{R}>\mathcal{T}$ then
14:   $\mathcal{F}_{C_{old}}=\mathcal{F}_{C_{old}}+1$
15:  end if
16:end for
17:return $\mathcal{F}$
The collision frequency for classes that do not exhibit collisions will be set to 1. To prevent the collision frequency of certain classes from becoming excessively large, we apply sigmoid smoothing. The results are then normalized to obtain $\mathcal{F^{\prime}}$, as in Eq. (11).
@@
TABLE I: mIoUmIoU on the AVSBench-CI dataset for different class-incremental audio-visual segmentation scenarios.

### v2

@@ Abstract
Recently, significant progress has been made in multi-modal continual learning, aiming to learn new tasks sequentially in multi-modal settings while preserving performance on previously learned ones. However, existing methods mainly focus on coarse-grained tasks, with limitations in addressing modality entanglement in fine-grained continual learning settings. To bridge this gap, we introduce a novel Continual Audio-Visual Segmentation (CAVS) task, aiming to continuously segment new classes guided by audio. Through comprehensive analysis, two critical challenges are identified: 1) multi-modal semantic drift, where a sounding objects is labeled as background in sequential tasks; 2) co-occurrence confusion, where frequent co-occurring classes tend to be confused. In this work, a Collision-based Multi-modal Rehearsal (CMR) framework is designed to address these challenges. Specifically, for multi-modal semantic drift, a Multi-modal Sample Selection (MSS) strategy is proposed to select samples with high modal consistency for rehearsal. Meanwhile, for co-occurence confusion, a Collision-based Sample Rehearsal (CSR) mechanism is designed, allowing for the increase of rehearsal sample frequency of those confusable classes during training process. Moreover, we construct three audio-visual incremental scenarios to verify effectiveness of our method. Comprehensive experiments demonstrate that our method significantly outperforms single-modal continual learning methods. We
1 Introduction
Humans are inherently capable of continuously learning while retaining knowledge from previous tasks. For example, infants can progressively recognize new animals while remembering those they have already learned. This human ability has motivated extensive research into continual learning Wang et al. (2024), which enables models to learn sequential tasks. Early work Rebuffi et al. (2017); Bang et al. (2021); Sun et al. (2023) primarily focused on classification, employing techniques such as regularization or rehearsal to mitigate catastrophic forgetting. Subsequent methods Cermelli et al. (2020) have extended continual learning to semantic segmentation. However, when directly applied to multi-modal (e.g. audio-visual) scenarios, these single-modal methods exhibit suboptimal performance Mo et al. (2023).
@@ 1 Introduction
Recently, several methods Mo et al. (2023); Pian et al. (2023; 2025); Yang et al. (2023) have extended continual learning to multi-modal scenarios. For example, AV-CIL Pian et al. (2023) proposes a continual audio-visual classification method with a dual similarity constraint enforcing both instance-level and class-level cross-modal semantic consistency. ContAV-Sep Pian et al. (2025) proposes a framework for audio-visual separation that incorporates cross-modal similarity distillation to preserve semantic consistency between modalities. Meanwhile, real-world applications require fine-grained audio-visual continual learning. For example, embodied intelligence needs to identify the source of a vocalization from environmental audio-visual cues. However, existing methods primarily focus on coarse-grained audio-visual tasks and therefore fail to address fine-grained tasks, such as disentangling pixel-level visual features from audio signals under continual learning scenarios.
Meanwhile, recent research Zhou et al. (2022; 2024); Yang et al. (2024) has explored fine-grained modality entanglement between audio signals and visual features in audio-visual segmentation. AVSBench Zhou et al. (2022) establishes the first benchmark for aligning the pixel-level visual semantics with the corresponding audio signals. COMBO Yang et al. (2024) further explores bilateral relations of three entanglements, pixel, modality, and temporal, to enhance the model’s representational capacity. However, audio-visual segmentation cannot be directly applied to continual learning scenarios, as it is designed for static settings.
To this end, we introduce a novel fine-grained multi-modal continual learning task, termed Continual Audio-Visual Segmentation (CAVS). Specifically, CAVS needs to perform audio-visual segmentation in a sequential task setting while retaining knowledge of previously seen classes. To address CAVS, we reformulate the AVS Zhou et al. (2022; 2024) framework and adapt classical continual semantic segmentation methods to the audio-visual context. Based on our observations, we identify two new challenges in fine-grained continual learning tasks: (1) Multi-modal semantic drift: Incorrect audio-visual semantic alignment (e.g. drum-background) due to mislabeling of learned classes as background exacerbates catastrophic forgetting. (2) Co-occurrence confusion: Frequent co-occurrence of categories leads to modality entanglement, for example, the audio modality of woman becomes entangled with visual modality of guitar in Fig. 1. In essence, these two issues are manifestations of modality entanglement from different perspectives.
@@
To validate the effectiveness of CMR, we reformulate the audio-visual dataset AVSBench Zhou et al. (2022) into three sequential task setup to better simulate a continual learning scenario. Specifically, our datasets include (1) AVSBench-Class Incremental (AVSBench-CI), (2) AVSBench-Class Incremental for Single-object (AVSBench-CIS), and (3) AVSBench-Class Incremental for Multi-object (AVSBench-CIM). Comprehensive experiments demonstrate that our proposed method achieves encouraging performance, showcasing its ability to effectively address the multi-modal semantic drift and co-occurrence confusion in CAVS.
@@
2 Related Works
2.1 Continual Learning.
Continual learning focuses on incrementally training models to adapt to new tasks while preserving knowledge from previously learned ones. Recently, many works  Rebuffi et al. (2017); Bang et al. (2021); Sun et al. (2023); ye2025domain; Li et al. (2024); Kang et al. (2022); Ostapenko et al. (2019); Shin et al. (2017); Wu et al. (2018); Chen et al. (2025) have proposed regularization-based and rehearsal-based methods to address the problem of catastrophic forgetting. Rehearsal-based methods Rebuffi et al. (2017); Bang et al. (2021); Sun et al. (2023); Li et al. (2024); Kang et al. (2022) allow for the storage of a small subset of old data in memory, which is later utilized for rehearsal during training. iCaRL Rebuffi et al. (2017) introduces a strategy to identify and retain the most representative samples for each class, which are replayed during training to mitigate forgetting in class-incremental learning. Pseudo-sample rehearsal-based methods Ostapenko et al. (2019); Shin et al. (2017); Wu et al. (2018) utilize generative models to create pseudo-samples of old classes. DGR Shin et al. (2017) establishes an initial framework where learning each new task is coupled with replaying the data generated by the old generative model. Building upon continual learning Rebuffi et al. (2017); Li et al. (2024); Kang et al. (2022); Ostapenko et al. (2019); Shin et al. (2017); Wu et al. (2018), Class-Incremental Semantic Segmentation (CISS) requires pixel-level classification to achieve fine-grained segmentation Cermelli et al. (2020); Douillard et al. (2021); Zhang et al. (2022); Yang et al. (2023); Cha et al. (2021); Yin et al. (2025). PLOP Douillard et al. (2021) suggests generating pseudo-labels by identifying latent past classes within the current background. ScaleSeg Yang et al. (2023) employs prototypes refined through online contrastive clustering and incorporates a background diversity strategy to boost plasticity. While  Cermelli et al. (2020) addressed semantic shifts within a single modality, our work reveals more complex multi-modal semantic drift where modal consistency is considered.
2.2 Audio Visual Segmentation
Audio-visual segmentation (AVS) is a novel and challenging task that localizes sound sources in visual scenes by pixel-level prediction Zhou et al. (2022; 2024); Yang et al. (2024); Mao et al. (2023); Li et al. (2023); Huang et al. (2025); Liu et al. (2025b; a); Wang et al. (2025). AVSBench Zhou et al. (2022) establishes the first audio-visual segmentation benchmark and introduces the Temporal Pixel-wise Audio-Visual Interaction (TPAVI) module to incorporate audio semantics as guidance for visual segmentation. AVSegFormer Zhou et al. (2024) develops a transformer-based framework with audio queries, learnable queries, and an audio-visual mixer for selective attention and dynamic feature adjustment. CATR  Li et al. (2023) proposes a combinatorial fusion framework that captures audio-visual spatiotemporal dependencies through cross-modal interaction modelling. ECMVAE Mao et al. (2023) decomposes audio and visual data in latent space, explicitly modeling both shared and modality-specific representations to enhance segmentation performance. COMBO Yang et al. (2024) rethinks AVS by exploring the bilateral relations of three entanglements, pixel, modality, and temporal, to enhance the model’s representation ability. In this work, we develop a framework for continual learning scenarios, making the task more aligned with real-world applications.
@@ 2.2 Audio Visual Segmentation
3 Methods
@@
3.1 Revisiting Continual Semantic Segmentation (CSS)
CSS assumes that tasks arrive sequentially, with each task containing a set of categories $\mathcal{C}^{t}$ and a corresponding training set $\mathcal{D}_{t}$, where $t$ denotes the current learning stage.
[…]
@@ 3.1 Revisiting Continual Semantic Segmentation (CSS)
3.2 Problem Setup and Notation of CAVS
For CAVS, the input space is defined as $\mathcal{S}\subset\mathcal{X}\times\mathcal{A}$, where $\mathcal{X}$ and $\mathcal{A}$ represent the visual and audio modalities, respectively. Each input sample $S=(\{S^{k}_{v}\},S_{a})\in\mathcal{D}_{t}$ contains $T$ consecutive video frames paired with an audio signal $S_{a}$, where $T=10$. The sounding objects in the $k$-th video frame $S^{k}_{v}$ are annotated with pixel-level labels. The objective of the $t$-th learning stage is to learn a model $f_{\theta^{t}}^{v,a}:\mathcal{S}\mapsto\mathbb{R}^{N\times\lvert{\mathcal{C}^{t}}\rvert}$, where $N$ is the number of pixels per frame. In this setting, the segmentation mask for pixel $i$ can be computed as follows:
In contrast, for task $\mathcal{D}_{t}$, both non-sounding objects from $\mathcal{Y}^{t}$ and sounding objects from $\mathcal{Y}^{t-1}$ are assigned the background label, while the audio $S_{a}$ remains unchanged. Compared to AV-ICL Pian et al. (2023), CAVS demands more substantial fine-grained alignment between global audio cues and local visual semantics.
Figure 3: Illustration of inter-modal consistent samples and entanglement of modality. (a) Cases 1 and 4 don’t appear in practice because selection uses already well-trained samples where audio and video predictions match the ground truth. Case 3 represents samples characterized by multi-modal semantic drift and is typically excluded due to substantially large $\lvert\Delta(S_{a})\rvert$. Conversely, Case 2 is kept because of its cross-modal semantic consistency. (b) Classes with infrequent co-occurrence exhibit weak audio-visual entanglement, while frequent co-occurrence leads to strong cross-modal entanglement (e.g., guitar sounds and images of women).
3.3 Multi-modal Sample Selection
@@
Inspired by the work in Wei et al. (2024a), where Shapley values are leveraged to quantify uni-modal contributions to model predictions, we propose a Multi-modal Sample Selection (MSS) strategy. By quantifying the contribution of the audio modality, this strategy identifies samples with high inter-modal consistency for rehearsal. Formally, given a video sample $S=(\{S^{k}_{v}\},S_{a})\in\mathcal{D}_{t}$, we train two parallel models:
$f_{\theta_{t}}^{v}(\{S^{k}_{v}\}):\mathcal{X}\mapsto\mathbb{R}^{N\times\lvert{\mathcal{Y}^{t}}\rvert}$, (1)
$f_{\theta_{t}}^{v,a}(\{S^{k}_{v}\},S_{a}):\mathcal{S}\mapsto\mathbb{R}^{N\times\lvert{\mathcal{Y}^{t}}\rvert}$. (2)
@@
For each newly added class $c\in\mathcal{C}^{t}$, we select the top-$k$ samples with the smallest absolute audio contribution deviation $\lvert\Delta(S_{a})\rvert$ from $\mathcal{D}_{t}$ to construct the memory buffer $M_{t}$.
These selected samples are stored and replayed during the training of subsequent tasks through $\mathcal{D}_{t+1}\cup M_{t}$, which effectively reinforces cross-modal associations. Our ablation studies demonstrate that this criterion outperforms random selection by 2.0 mIoU (see Tab. 3), highlighting the importance of modality consistency in sample rehearsal.
3.4 Collision-based Sample Rehearsal
@@
To implement this idea, we propose the Collision-based Sample Rehearsal (CSR) strategy, which identifies classes prone to co-occurrence confusion by detecting collisions between the old model’s predictions and the ground truth. As illustrated in Fig. 4, for a new sample $S$, a collision occurs when the old model $f^{v,a}_{\theta_{t-1}}$ predicts an old class $c_{old}\in\mathcal{Y}^{t-1}$ in a spatial position where the ground truth $c_{new}\in\mathcal{C}^{t}$ appears.
Algorithm 1 Collision-Based Sampling
1:Old model $f^{v,a}_{\theta^{t-1}}$, Training dataset $\mathcal{D}_{t}$, Semantic label $\mathcal{Y}_{\mathrm{gt}}$, Threshold $\mathcal{T}$
2:Collision frequency $\mathcal{F}$
3:for $S_{i}\in\mathcal{D}_{t}$ do
4:  Compute $\hat{\mathcal{Y}}\leftarrow f^{v,a}_{\theta^{t-1}}(S_{i})$
5:  Mask $\mathcal{M}\leftarrow(\hat{\mathcal{Y}}\neq\text{background})\land(\mathcal{Y}\neq\text{background})$
6:  Collision Region $\mathcal{I}\leftarrow(\hat{\mathcal{Y}}\neq\mathcal{Y})\land\mathcal{M}$
[…]
8:  for $i\in\mathcal{I}$ do
9:   $\mathrm{Collision}(\hat{\mathcal{Y}}_{i},\mathcal{Y}_{i})\leftarrow\mathrm{Collision}(\hat{\mathcal{Y}}_{i},\mathcal{Y}_{i})+1$
10:  end for
[…]
12:  $(\mathcal{C}_{\mathrm{old}},\mathcal{C}_{\mathrm{new}})\leftarrow\arg\max\mathrm{Collision}(\hat{\mathcal{Y}}_{\mathcal{I}},\mathcal{Y}_{\mathcal{I}})$
13:  $\mathcal{R}\leftarrow\dfrac{\mathrm{Collision}(\mathcal{C}_{\mathrm{old}},\mathcal{C}_{\mathrm{new}})}{\sum\mathrm{Collision}(\hat{\mathcal{Y}}_{\mathcal{I}},\mathcal{Y}_{\mathcal{I}})}$
[…]
15:  if $\mathcal{R}>\mathcal{T}$ then
16:   $\mathcal{F}_{\mathcal{C}_{\mathrm{old}}}\leftarrow\mathcal{F}_{\mathcal{C}_{\mathrm{old}}}+1$
17:  end if
18:end for
19:return $\mathcal{F}$
Specifically, with the old model and task $\mathcal{D}_{t}$, the collisions between the prediction of $f^{v,a}_{\theta_{t-1}}$ and $\mathcal{D}_{t}$ is first computed. Inferring the video $S$ with the old model $f^{v,a}_{\theta_{t-1}}$, we obtain a collision pair $(c_{old},c_{new})$. Since the old model has not trained on new samples, it can only predict old classes $c_{old}\subset\mathcal{Y}^{t-1}$. Assuming that the predicted result is $c_{old}$ and the ground truth label is $c_{gt}$, we count all collision pairs $(c_{old},c_{new})$ and identify the learned class with the highest number of collisions as the most confusing class for the current video $S$:
$\mathcal{P}(S)=\arg\max\{Count(c_{i},c_{j})\lvert i\in\mathcal{Y}^{t-1},j\in\mathcal{C}^{t}\}.$ (6)
@@
$R_{c}=\frac{Count(\mathcal{P}(S)=c)}{\sum\{Count(c_{i},c_{j})\lvert i\in\mathcal{Y}^{t-1},j\in\mathcal{C}^{t}\}},$ (7)
@@
The collision frequency for classes that do not exhibit collisions will be set to 1. To prevent the collision frequency of certain classes from becoming excessively large, we apply sigmoid smoothing. The results are then normalized to obtain $\mathcal{F^{\prime}}$, as in Eq. equation 9.
@@
4 Experiments
4.1 AVSBench Datasets
In our work, a class-incremental audio-visual segmentation dataset (AVSBench-CI) is constructed from the well-known dataset AVSBench-semantic Zhou et al. (2024) to validate the proposed CMR. AVSBench-semantic utilizes the techniques introduced in VGGSound Chen et al. (2020) to collect videos, ensuring that the audio and visual clips align with the intended semantics. The dataset provides semantic segmentation maps for videos as labels to enhance audio-visual semantic segmentation (AVSS). It contains a total of 11,356 videos spanning 70 categories. Each video segment consists of 10 frames of images and one 10-second audio clip. We divide the 70 categories in AVSBench-semantic for the original dataset into three training steps: 60-10, 60-5, and 65-1. Following the conventional continual semantic segmentation setup, the three training steps are divided into overlapped and disjoint settings to evaluate the model’s performance under different task stream configurations.
In the overlapped setting, the classes are divided sequentially, meaning that classes from past and future tasks may appear in the current data and be labelled as background. In the disjoint setting, a community detection algorithm Sahu (2024) is employed to minimize the overlap of training data between consecutive steps. Therefore, the current data will not contain classes from future or past tasks. This setup closely aligns with continual learning scenarios. Furthermore, we expand the single-semantic dataset (AVSBench-CIS) and the multi-semantic dataset (AVSBench-CIM) based on the number of targets in the videos. AVSBench-CIS and AVSBench-CIM address scenarios involving modality entanglement with single-target and multi-target settings, respectively. The same settings are applied to these datasets.
4.2 Experimental Setup
4.2.1 Baselines
Since semantic segmentation can be regarded as a pixel-wise classification task, we compare our method with both classification and segmentation methods. Baselines, evaluation metrics, and implementation details are provided in the appendix.
Table 1: mIoUmIoU on the AVSBench-CI dataset for different class-incremental audio-visual segmentation scenarios.